Find maximum value of y(x) = integral from 0 to x (t+sint).dt in [-1,1]

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
y(x)=integration(t+sint)d=[(t^2)/2 -cos(t)](between 0 to x)=(x^2)/2 -cos(x)+1
max of function will occur when cox(x) is minimum, y(x)max =1/2-cos(1)+1=3/2 -cos(1)

Thanks & Regards
Sher Mohammad,
askIITians Faculty
B.tech
IIT Delhi.